OpenStudy (anonymous):

SImplify.. 4/x*(6/x+4-6/x-4)

OpenStudy (goformit100):

@nincompoop

OpenStudy (anonymous):

\[\frac{ 4 }{ x }*(\frac{ 6 }{ x+4 }-\frac{ 6 }{ x-4 })\] is this the prob?

OpenStudy (anonymous):

@ gabby, is that the prob?

OpenStudy (anonymous):

yes sorry